Research Institute for Mathematical Sciences Workshop

Analysis on homogeneous spaces and representation of Lie groups

August 4 – August 8, 1997

Conference Hall

Yukawa Institute for Theoretical Physics (the next door of RIMS)

Program

August 4 (MON):

13:30 – 14:20 : Masaki Kashiwara (RIMS)

Integral transform of D-modules and representation theory

14:40 – 15:30 : Kari Vilonen (Brandeis Univ.)

Geometric realization of Whittaker functions and the Langlands conjecture

15:50 – 16:40 : Toshiyuki Tanisaki (Hiroshima Univ.)

Hypergeometric systems and Radon transforms for Hermitian symmetric spaces

August 5 (TUE):

10:00 – 10:50 : Wilfried Schmid (Harvard Univ.)

TBA

11:10 – 12:00 : Takahiro Hayata (Univ. Tokyo)

Shintani-Whittaker functions on real semisimple Lie groups of low rank

13:30 – 14:20 : David Vogan (MIT)

A Langlands classification for unitary representations

14:40 – 15:30 : Kaoru Hiraga (Kyoto Univ.)

On the multiplicity of the discrete series representations of semisimple groups

15:50 – 16:40 : Zhu Chengbo (National Univ. of Singapore)

On the cohomology of certain unitary representations August 6 (WED):

10:00 – 10:50 : Masato Wakayama (Kyushu Univ.)

Harmonic oscillators with matrix-valued symbols

11:10 – 12:00 : Erik Balslev (Aarhus Univ.)

Perturbation theory for Laplacians on hyperbolic Riemann surfaces

13:30 – 14:20 : Adam Korányi (Lehman College)

Equivariant differential operators on vector bundles over G/MAN

14:40 – 15:30 : George Tomanov (Univ. Claude Bernard Lyon I)

Group actions on homogeneous spaces and Oppenheim’s conjecture for unitary forms over skew fields

15:50 – 16:40 : Yves Benoist (Univ. Paris 7)

Linear groups with positive eigenvalues

August 7 (THU):

10:00 – 10:50 : Minoru Wakimoto (Kyushu Univ.)

Modular transformation of twisted admissible characters

11:10 – 12:00 : Fumihiro Sato (Rikkyo Univ.)

Weakly spherical homogeneous spaces and zeta functions

13:30 – 14:20 : Bertram Kostant (MIT)

The Toda Lattice and the quantum cohomology of the flag manifold

14:40 – 15:30 : Toru Umeda (Kyoto Univ.)

On Capelli Identities

15:50 – 16:40 : Tuong Ton-That (Univ. Iowa)

Poincaré-Birkhoff-Witt theorems and generalized Casimir Invariants for some infinite-dimensional Lie Groups

August 8 (FRI):

10:00 – 10:50 : Hiroyuki Ochiai (Rikkyo Univ.)

Spherical functions for an integrable system of rank two

11:10 – 12:00 : Ivan Cherednik (Univ. North Carolina at Chapel Hill)

Harish-Chandra transform and difference operators